# Disentangling Recognition and Decision Regrets in Image-Based Reinforcement Learning

### v1

@@ 5.1 Examples in Maze Environments
We consider simple mazes given on a $3$-by-$3$ grid, where the agent needs to navigate to some goal square on the grid. We consider different maze configurations, but in each configuration, the goal is separated from the agent by a locked door, and to be able to reach the goal, the agent needs to first pick up a key and unlock the door. With uniform probability, the door can either be at (i) the North-end, (ii) the Center, or (iii) the South-end of the parting wall between the goal and the agent (see Figure 4).
@@
We construct different agents by applying different pre-filters to input images as part of the recognition policy before feeding them forward to the same network architecture. This technique allows us to freely control the properties of an agent’s recognition policy and illustrate cases of over-specific vs. under-specific representations. We primarily consider three agents/pre-filters (see the corresponding column in Table 1, results for additional agents can be found in the appendix): (i) Identity makes no modifications to the input image. This means the key-door color is preserved as an irrelevant feature, making Identity susceptible to learning over-specific representations. (ii) HideDoor hides the location of the true door by placing false doors in the maze but in the same color as the original door. This means HideDoor can only rely on the key-door color as a spurious feature to navigate during training, which may lead to learning under-specific representations. (iii) HideColors re-colors the key and the door as gray, the same color as the keys and the door in the test environment. This should make HideColors easily generalizable to the test environment.
@@ 5.2 Examples in the Atari Game Pong
Setup
Agents
Similar to before, we construct different agents via different pre-filters: (i) Identity makes no modifications to the input image. (ii) NoCounter hides the frame counter keeping the distractions. This should leave the agent susceptible to learning over-specific representations that include irrelevant features extracted from distractions. (iii) JustCounter hides everything but the game counter. This should leave the agent susceptible to learning under-specific representations that only include a spurious feature, which is the frame count. (iv) JustGameField hides both the frame counter (spurious) and any distractions (irrelevant), which is ideal.
@@ Appendix A A Worked Example for Generalization
${G\!R}$ $={\max}{}_{\rho,\pi}V_{\varepsilon^{\text{test}}}(\rho,\pi)-V_{\varepsilon^{\text{test}}}(\operatorname{argmax}{}_{\pi}V_{\varepsilon^{\text{train}}}(\rho_{0},\pi))$ (13)
$=4.737-2.360=2.377.$ (14)
@@
$G\!R{}^{\text{rec}}$ $=V_{\varepsilon^{\text{test}}}(\operatorname{argmax}{}_{\rho,\pi}V_{\varepsilon^{\text{test}}}(\rho,\pi))-V_{\varepsilon^{\text{test}}}(\rho_{0},\operatorname{argmax}{}_{\pi}V_{\varepsilon^{\text{test}}}(\rho_{0},\pi))$ (15)
$=4.737-2.989=1.748$ (16)
$G\!R{}^{\text{dec}}$ $=V_{\varepsilon^{\text{test}}}(\rho_{0},\operatorname{argmax}{}_{\pi}V_{\varepsilon^{\text{test}}}(\rho_{0},\pi))-V_{\varepsilon^{\text{test}}}(\rho_{0},\operatorname{argmax}{}_{\pi}V_{\varepsilon^{\text{train}}}(\rho_{0},\pi))$ (17)
$=2.989-2.360=0.629$ (18)
@@
We now consider a different recognition policy $\rho_{1}$ that focuses on the timestamp in the top-right corner of each image in Fig. 2. In particular the representation state is defined by $\rho^{\prime}_{1}(x_{t})=0$ if $t=0$, otherwise $\rho^{\prime}_{1}(x_{t})=1$ if $t$ is odd and $\rho^{\prime}_{1}(x_{t})=2$ if $t$ is even. This allows us to obtain the optimal reward in $\varepsilon^{\text{train}}$ since if we apply the optimal decision policy (that takes action $1$ in states $0,1$ and action 00 in state $1$), then we will follow the optimal trajectory and $s_{t}=\rho^{\prime}_{1}(x_{t})$ at all times $t$. Note that in this optimal trajectory the action sequence is $1,1,0,1,0,1,\ldots$
@@
${G\!R}$ $={\max}{}_{\rho,\pi}V_{\varepsilon^{\text{test}}}(\rho,\pi)-V_{\varepsilon^{\text{test}}}(\operatorname{argmax}{}_{\pi}V_{\varepsilon^{\text{train}}}(\rho_{0},\pi))$ (19)
$=4.737-0.900=3.837.$ (20)
@@
$G\!R{}^{\text{rec}}$ $=V_{\varepsilon^{\text{test}}}(\operatorname{argmax}{}_{\rho,\pi}V_{\varepsilon^{\text{test}}}(\rho,\pi))-V_{\varepsilon^{\text{test}}}(\rho_{1},\operatorname{argmax}{}_{\pi}V_{\varepsilon^{\text{test}}}(\rho_{1},\pi))$ (21)
$=4.737-4.737=0.$ (22)
$G\!R{}^{\text{dec}}$ $=V_{\varepsilon^{\text{test}}}(\rho_{1},\operatorname{argmax}{}_{\pi}V_{\varepsilon^{\text{test}}}(\rho_{1},\pi))-V_{\varepsilon^{\text{test}}}(\rho_{1},\operatorname{argmax}{}_{\pi}V_{\varepsilon^{\text{train}}}(\rho_{1},\pi))$ (23)
$=4.737-0.900=3.837.$ (24)
@@
Appendix B Extended Results
@@
Appendix C Detailed Description of the Experiments
@@
We consider the same policy architecture as in Appendix C.1 given for Minigrid. However this time, we consider the custom maze configurations given in 4. Agents are trained in $64$ parallel environments for $50,\!000$ total time steps. We use the default hyper-parameters for the PPO algorithm in the skrl package [^3] The number of learning epochs during each update is $8$, the number of minibatches during each learning epoch is $2$, the discount factor ($\gamma$) is $0.99$, the coefficient $\lambda$ used for generalized advantage estimation is $0.95$, the learning rate is $0.001$ (an Adam optimizer is used), the clipping coefficient for the norm of the gradients is $0.5$, and the clipping coefficient for computing the clipped surrogate objective is $0.2$. , except we collect $4096$ total transitions in between each policy update, set the entropy loss scale as $0.01$ and the value loss scale as $0.5$, and activate the running standard scalers for output values. Agents are evaluated by rolling out $1024$ full episodes. We repeat all our experiments five times to obtain 1-sigma error bars.

### v2

@@ 5.1 Examples in Maze Environments
We consider simple mazes given on a $3$-by-$3$ grid, where the agent needs to navigate to some goal square. We consider different maze configurations, but in each configuration, the goal is separated from the agent by a locked door, and to be able to reach the goal, the agent needs to first pick up a key and unlock the door. With uniform probability, the door can either be at (i) the North-end, (ii) the Center, or (iii) the South-end of the parting wall between the goal and the agent (see Figure 4).
@@
We construct different agents by applying different pre-filters to input images as part of the recognition policy before feeding them forward to the same network architecture. This technique allows us to freely control the properties of an agent’s recognition policy and illustrate cases of over-specific vs. under-specific representations. We primarily consider three agents/pre-filters (see the corresponding column in Table 1, results for additional agents can be found in the supplementary material): (i) Identity makes no modifications to the input image. This means the key-door color is preserved as an irrelevant feature, making Identity susceptible to learning over-specific representations. (ii) HideDoor hides the location of the true door by placing false doors in the maze but in the same color as the original door. This means HideDoor can only rely on the key-door color as a spurious feature to navigate during training, which may lead to learning under-specific representations. (iii) HideColors re-colors the key and the door as gray, the same color as the keys and the door in the test environment. This should make HideColors easily generalizable to the test environment.
@@ 5.2 Examples in the Atari Game Pong
Setup
Agents
Similar to before, we construct different agents via different pre-filters (see the corresponding column in Table 2): (i) Identity makes no modifications to the input image. (ii) NoCounter hides the frame counter keeping the distractions. This should leave the agent susceptible to learning over-specific representations that include irrelevant features extracted from distractions. (iii) JustCounter hides everything but the game counter. This should leave the agent susceptible to learning under-specific representations that only include a spurious feature, the frame count. (iv) JustGameField hides both the frame counter (spurious) and any distractions (irrelevant), which is ideal.
@@ Results
5.3 Use Case: Model Selection
Figure 5: Regret and Generalization Regrets of Identity in Maze Environments. Reducing the number of latent channels prevents over-specific representations and lowers $\widehat{G\!R}{}^{\text{dec}}$. However, this also results in the recognition policy overfitting to the key-door color, ultimately leading to no improvements in terms of the overall generalization regret.
Analyzing the recognition and decision generalization regrets of RL models can help make informed decisions about how to improve generalization. If generalization performance is poor due to over-specific representations, we may switch to an algorithm that is capable of learning invariances within the observation space (e.g. Sonar et al., 2021). Conversely, if our representations are under-specific, we may introduce additional loss terms to encourage the model to extract task-critical features.
As a concrete example, we consider one of the simplest adjustments to make to the model architecture: changing the representation dimensionality. We consider Identity from the examples in maze environments, whose recognition policy outputs a latent representation with $642^{6}=$ 64 channels. Observing that Identity has a high $\widehat{G\!R}{}^{\text{dec}}$ relative to $\widehat{G\!R}{}^{\text{rec}}$, we conclude that generalization fails due to over-specific representations. To discourage the model from extracting irrelevant features, we decide to reduce the number of latent channels. We can see in Figure 5 that our intervention is successful in lowering $\widehat{G\!R}{}^{\text{dec}}$, but the model then ends up overfitting to spurious features like the key-door color instead, resulting in high $\widehat{G\!R}{}^{\text{rec}}$ and no overall improvement in terms of the generalization regret. Now that we know our setting suffers both from irrelevant and spurious features, we can start investigating alternative ways to address either one of these problematic features, adopting an appropriate number of latent channels to the issue we target. For instance, we can also see in Figure 5 that the increase in $\widehat{G\!R}{}^{\text{rec}}$ is rather marginal until the number of latent channels is reduced to $162^{4}=$ 16. We can set this as our new representation size, knowing it is sufficiently large to capture all task-critical information, and decide to address the issue of irrelevant features first.
6 Conclusion
In this paper, we introduced the notions of recognition and decision regrets, which we used to analyze the generalization performance of RL agents. Through illustrative examples, we demonstrated how these regret metrics can characterize under-specific and over-specific representations, offering a more nuanced perspective on observational overfitting in image-based RL. We highlighted how such characterization could be helpful in systematically improving the generalization performance of RL models. Future research should focus on identifying specific strategies that one can follow when an RL model fails to generalize due to over-specific vs. under-specific representations. Our code is publicly available at https://github.com/alihanhyk/disentangling-regret.
@@ Appendix A A Worked Example for Generalization
${G\!R}={\max}{}_{\rho,\pi}V_{\varepsilon^{\text{test}}}(\rho,\pi)-V_{\varepsilon^{\text{test}}}(\operatorname{argmax}{}_{\pi}V_{\varepsilon^{\text{train}}}(\rho_{0},\pi))=4.737-2.360=2.377$ (13)
@@
$G\!R{}^{\text{rec}}$ $=V_{\varepsilon^{\text{test}}}(\operatorname{argmax}{}_{\rho,\pi}V_{\varepsilon^{\text{test}}}(\rho,\pi))-V_{\varepsilon^{\text{test}}}(\rho_{0},\operatorname{argmax}{}_{\pi}V_{\varepsilon^{\text{test}}}(\rho_{0},\pi))$
$=4.737-2.989$
$=1.748$ (14)
$G\!R{}^{\text{dec}}$ $=V_{\varepsilon^{\text{test}}}(\rho_{0},\operatorname{argmax}{}_{\pi}V_{\varepsilon^{\text{test}}}(\rho_{0},\pi))-V_{\varepsilon^{\text{test}}}(\rho_{0},\operatorname{argmax}{}_{\pi}V_{\varepsilon^{\text{train}}}(\rho_{0},\pi))$
$=2.989-2.360$
$=0.629$ (15)
@@
We now consider a different recognition policy $\rho_{1}$ that focuses on the timestamp in the top-right corner of each image in Figure 2. In particular the representation state is defined by $\rho^{\prime}_{1}(x_{t})=0$ if $t=0$, otherwise $\rho^{\prime}_{1}(x_{t})=1$ if $t$ is odd and $\rho^{\prime}_{1}(x_{t})=2$ if $t$ is even. This allows us to obtain the optimal reward in $\varepsilon^{\text{train}}$ since if we apply the optimal decision policy (that takes action $1$ in states $0,1$ and action 00 in state $1$), then we will follow the optimal trajectory and $s_{t}=\rho^{\prime}_{1}(x_{t})$ at all times $t$. Note that in this optimal trajectory the action sequence is $1,1,0,1,0,1,\ldots$
@@
${G\!R}$ $={\max}{}_{\rho,\pi}V_{\varepsilon^{\text{test}}}(\rho,\pi)-V_{\varepsilon^{\text{test}}}(\operatorname{argmax}{}_{\pi}V_{\varepsilon^{\text{train}}}(\rho_{0},\pi))=4.737-0.900=3.837$ (16)
@@
$G\!R{}^{\text{rec}}$ $=V_{\varepsilon^{\text{test}}}(\operatorname{argmax}{}_{\rho,\pi}V_{\varepsilon^{\text{test}}}(\rho,\pi))-V_{\varepsilon^{\text{test}}}(\rho_{1},\operatorname{argmax}{}_{\pi}V_{\varepsilon^{\text{test}}}(\rho_{1},\pi))$
$=4.737-4.737$
$=0$ (17)
$G\!R{}^{\text{dec}}$ $=V_{\varepsilon^{\text{test}}}(\rho_{1},\operatorname{argmax}{}_{\pi}V_{\varepsilon^{\text{test}}}(\rho_{1},\pi))-V_{\varepsilon^{\text{test}}}(\rho_{1},\operatorname{argmax}{}_{\pi}V_{\varepsilon^{\text{train}}}(\rho_{1},\pi))$
$=4.737-0.900$
$=3.837$ (18)
@@
Appendix B Extended Results for Maze Environments
@@
Appendix C Details of the Experiments
@@
We consider the same policy architecture as in Appendix C.1 given for Minigrid. However this time, we consider the custom maze configurations given in Figure 4. Agents are trained in $64$ parallel environments for $50,\!000$ total time steps. We use the default hyper-parameters for the PPO algorithm in the skrl package [^2] The number of learning epochs during each update is $8$, the number of minibatches during each learning epoch is $2$, the discount factor ($\gamma$) is $0.99$, the coefficient $\lambda$ used for generalized advantage estimation is $0.95$, the learning rate is $0.001$ (an Adam optimizer is used), the clipping coefficient for the norm of the gradients is $0.5$, and the clipping coefficient for computing the clipped surrogate objective is $0.2$. , except we collect $4096$ total transitions in between each policy update, set the entropy loss scale as $0.01$ and the value loss scale as $0.5$, and activate the running standard scalers for output values. Agents are evaluated by rolling out $1024$ full episodes. We repeat all our experiments five times to obtain 1-sigma error bars.
@@
C.4 Experiments in Section 5.3
We consider the same setup as in Section 5.1. Only difference is that we vary the channel size of the last convolutional layer. During these experiments, especially for models with very small channel sizes, PPO tends to output unstable solutions. To account for the increased variation, we repeat all our experiments ten times instead of five to obtain 1-sigma error bars.